Patrons on an amusement park ride called the Rotor stand with their backs against the wall of a rotating cylinder while the floor drops away beneath them. To keep from sliding downward, they require a centripetal acceleration in excess of about 25 ms^-2. The rotor has a diameter of 5.0 m. What is the minimum frequency for its rotation?

1
Expert's answer
2021-10-31T18:11:52-0400

Solution.

"a=25 m\/s^2;"

"d=5m;"

"n -?;"

"a=\dfrac{v^2}{r}\implies v=\sqrt{ar};"

"v=\sqrt{25\sdot2.5}=\sqrt{62.5}=7.9 m/s;"

"v=2\pi rn\implies n=\dfrac{v}{2\pi r};"

"n=\dfrac{7.9}{2\sdot 3.14\sdot2.5}=0.5s^{-1};"